OpenStudy (anonymous):

Solve, using linear combination. 3x – y = 4 2x + y = 6

OpenStudy (haseeb96):

3x-y=4...............eq#1 2x+y=6...............eq#2 take eq#1 y=3x-4...........eq#A put Eq#A into Eq#2 2x+3x-4=6 5x=6+4 5x=10 x=2 now put x=2 into eq#A y=3(2)-4 y=6-4 y=2 so the solution set is { (2,2) } this is the answer

OpenStudy (anonymous):

thankyou!! can you help me with another one!

OpenStudy (anonymous):

Use the graph to state the solution for the system. x – y = 3 (line a) x + y = 3 (line b)

OpenStudy (anonymous):